OpenStudy (anonymous):

the empire state building is 113 feet more than three times the height of the hyperion. the empire state building is 1250 ft. how tall is the hyperion and write an equation

OpenStudy (anonymous):

Let e be the empire state building's height and h be the hyperion's height. You can make two equations from the given information. \[e = 113+3h\]and \[e = 1250\]Now you can substitute e in the first equation, which will eliminate a variable (e) and allow you to solve for the 2nd variable (h)\[1250 = 113+3h\]Subtract 113 from both sides\[1137 = 3h\]Divide both sides by 3\[379 = h\]